40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 11 04:32:53 2010, yuta, Summary, CDS, found poor contact of DAC cable, previous A2L results were wrong 
    Reply  Thu Nov 11 11:51:30 2010, Koji, Summary, CDS, found poor contact of DAC cable, previous A2L results were wrong 
       Reply  Thu Nov 11 23:35:23 2010, Koji, Summary, CDS, found poor contact of DAC cable, previous A2L results were wrong IMG_3698.jpgbad-boys.pngIMG_3705.jpgIMG_3706.jpg
Message ID: 3901     Entry time: Thu Nov 11 23:35:23 2010     In reply to: 3895
Author: Koji 
Type: Summary 
Category: CDS 
Subject: found poor contact of DAC cable, previous A2L results were wrong 

[Koji / Yuta]

There were the guys who used the PENTEK 40pin connectors into the IDC 40pin connectors.
Those connectors are not compatible at all.

==> We replaced the connectors on the cables from DAC to IDC adapters to the dewhitening board for the vertex SUSs.

In addition, I found one of the Binary OUT IDC50pin connector has no clamp.

==> We put the IDC50pin clamp on it.

bad-boys.png

PENTEK connectors were inserted. The latches are not working!

IMG_3698.jpg

the vertical pitch is different between PENTEK and IDC!
IMG_3705.jpg

Wow! Where is the clamp???

IMG_3706.jpg

Quote:

The cause is apparent! The connectors on the cables are wrong!
Currently only 50% of the pin length goes into the connector!

Quote:

(Koji, Jenne, Yuta)

We found one of DAC cables had a poor contact.
That probably caused our too much "tilt" of the beam into MC.

  It was because one of the DAC cables(labeled CAB_1Y4_88) had a poor contact.
  If I push it really hard, it is ok. But maybe we'd better replace the cable.

What caused a poor connection?:
  I don't know.
  A month ago, we checked that they are connected, but things change.

 

 

ELOG V3.1.3-